A solid metallic sphere of diameter 28 cm is melted and number of cones formed by the diameter 4 2/3 cm and height 3 cm .find the no. Of cones formed .......

Answer by addingup(3677)

Radius = diameter/2 = 28/2 = 14
Volume of the sphere:
4/3(Pi*r^3) = Volume
4/3(Pi*14^3) = 11,494 This is how much material you have to form cones
-------------------------------------
Volume of a cone:
1/3(B*h) B: base and h: height. Since the base is Pi*r^2, we can re-write:
1/3((Pi*r^2)h)
1/3((Pi*14^2)3) = 615.75
---------------------------------------
11,494/615.75 = 18.667
You can make 18 cones and have some left over, 0.667 which is enough for 2/3 of a cone.